OpenStudy (anonymous):

here is a easier one simplify (√3+√10)(√3-10)

OpenStudy (anonymous):

last one to get it is a rotten egg

OpenStudy (anonymous):

is it \[(\sqrt{3}+\sqrt{10})(\sqrt{3}-\sqrt{10})\]?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

my bad forgot a root part

OpenStudy (anonymous):

\[(a+b)(a-b)=a^2-b^2\] so 3-10

OpenStudy (anonymous):

no its √3- √10

OpenStudy (anonymous):

possible answers 7√3, 13, -7√10 , or -7